# Transforming Indoor Localization: Advanced Transformer Architecture for NLOS Dominated Wireless Environments with Distributed Sensors

### v1

@@ Abstract
Indoor localization in challenging non-line-of-sight (NLOS) environments often leads to mediocre accuracy with traditional approaches. Deep learning (DL) has been applied to tackle these challenges; however, many DL approaches overlook computational complexity, especially for floating-point operations (FLOPs), making them unsuitable for resource-limited devices. Transformer-based models have achieved remarkable success in natural language processing (NLP) and computer vision (CV) tasks, motivating their use in wireless applications. However, their use in indoor localization remains nascent, and directly applying Transformers for indoor localization can be both computationally intensive and exhibit limitations in accuracy. To address these challenges, in this work, we introduce a novel tokenization approach, referred to as Sensor Snapshot Tokenization (SST), which preserves variable-specific representations of power delay profile (PDP) and enhances attention mechanisms by effectively capturing multi-variate correlation. Complementing this, we propose a lightweight Swish-Gated Linear Unit-based Transformer (L-SwiGLU Transformer) model, designed to reduce computational complexity without compromising localization accuracy. Together, these contributions mitigate the computational burden and dependency on large datasets, making Transformer models more efficient and suitable for resource-constrained scenarios. The proposed tokenization method enables the Vanilla Transformer to achieve a 90th percentile positioning error of 0.388 m in a highly NLOS indoor factory, surpassing conventional tokenization methods. The L-SwiGLU ViT further reduces the error to 0.355 m, achieving an 8.51% improvement. Additionally, the proposed model outperforms a 14.1 times larger model with a 46.13% improvement, underscoring its computational efficiency.
@@ I Introduction
The rapid advancement of wireless communication technologies, fueled by the deployment of 5G and the emerging vision of 6G, is driving transformative innovations in location-based services. Applications such as vehicle-to-everything (V2X) communication, urban air mobility (UAM), extended reality (XR), near-real-time robotic operations, and Smart X systems [1, 2] are at the forefront of this technological evolution. These advancements require the large-scale integration of Internet of Things (IoT) infrastructures, thereby intensifying the demand for accurate and efficient localization methods with reduced computational complexity to ensure the stringent quality of service (QoS) necessary for these cutting-edge applications. Effective indoor localization is particularly critical for applications such as navigation, safety and rescue operations, and resource optimization for customized user services. However, the need for reduced computational complexity is paramount, as most indoor positioning devices are constrained by limited computing power. Addressing these challenges requires the development of efficient and precise localization methods capable of operating within these constraints while supporting the demands of modern wireless communication systems.
@@
Global Navigation Satellite Systems (GNSS), commonly used for outdoor localization, face significant challenges in indoor environments due to severe channel conditions such as shadowing, fading, and noise, as well as the high probability of non-line-of-sight (NLOS) situations. Conventional wireless signal-based localization techniques [3], including Time of Arrival (TOA), Time Difference of Arrival (TDOA), Time of Flight (TOF), and Angle of Arrival (AOA), often perform poorly in highly NLOS environments, which are typical in indoor settings like factories, hospitals, shopping malls, and military installations. Additionally, these techniques frequently rely on empirical models that exhibit high complexity and a strong dependence on varying channel conditions.
@@
Fingerprinting [4] has emerged as a widely adopted approach for indoor localization, relying on signal measurements such as Received Signal Strength Indicator (RSSI) or Channel State Information (CSI) to create location-specific fingerprints. However, traditional machine learning (ML) based fingerprinting methods often struggle to adapt to environmental variability and fail to capture the complex spatial and temporal relationships in the data. To address these limitations, researchers have increasingly turned to deep learning (DL) techniques, such as Convolutional Neural Networks (CNNs) and Multi-Layer Perceptrons (MLPs) [3], which are capable of learning more robust and discriminative features from the fingerprint data. DL-based fingerprinting techniques have shown promising results in various studies, but such techniques also increase system complexity, requiring substantial memory and processing power. This added complexity can hinder real-time implementation and scalability. Therefore, addressing these challenges is crucial, and this paper aims to present an approach that balances performance improvements with reduced system overhead, making fingerprinting techniques more efficient and adaptable.
@@
The Transformer model [5] offers several advantages over traditional sequence processing techniques such as Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks, including parallelization capability, reduced training time, and improved handling of long-range dependencies [6]. Transformers, despite their proven success in fields such as natural language processing (NLP), computer vision (CV), and machine translation [5, 7], have not been extensively used for indoor localization. This underutilization is largely due to the lack of inductive biases such as locality and translation invariance, which are naturally present in CNNs but absent in Transformers, and due to the Transformer’s dependence on large datasets to capture both local and global features [7]. These challenges are further amplified by the way inputs are tokenized, which can impact the model’s training effectiveness.
@@
In highly NLOS indoor wireless environments, multi-antenna systems face significant challenges due to multipath propagation, where signals undergo reflection, diffraction, and scattering from walls, furniture, and other obstacles. This complex propagation leads to signal superposition, causing phase shifts, fading, and interference, which further complicate the resolution of spatial and temporal variations in such environments. Moreover, in the context of indoor localization, access to large-scale datasets is limited, as collecting data of this magnitude is both computationally expensive and cost-inefficient. Additionally, existing network infrastructures lack the high computational resources necessary to test models with billions of FLOPs, which further constrains the deployment of computationally intensive models. These challenges underscore the need for more efficient approaches that reduce the dependence of Transformer-based models on large datasets while simultaneously lowering their computational complexity, enabling practical deployment in complex indoor wireless environments.
@@
One critical yet often overlooked aspect of Transformers, particularly within the wireless communication domain, is tokenization. Conventional tokenization methods typically aggregate multiple variables, such as delayed events and physical measurements, into a single representation. This aggregation can obscure variable-specific features, leading to less informative attention maps and degrading model performance. Consequently, the model’s dependency on extensive datasets increases, as larger datasets are required to effectively map complex multi-antenna radio signals to precise locations. Therefore, in this paper, we investigate the Transformer architecture and enhance its learning capabilities by proposing a sophisticated and physically interpretable tokenization technique referred to as Sensor Snapshot Tokenization (SST) that leverages the characteristics of wireless communication systems, specifically addressing inherent channel independence and enabling the Transformer to effectively learn the multivariate correlations using the tokens form multi-antenna radio system. This approach facilitates the development of variate-centric representations, resulting in meaningful attention maps and leading to a decreased dependence of the Transformer on large models and extensive datasets.
To further enhance efficiency and reduce computational complexity, we introduced modifications to several components of the vanilla Transformer architecture. These include replacing conventional normalization layers, redesigning the feed-forward network, and redefining the prediction mechanism using tokens from the final encoder block. The resulting enhanced Transformer model is termed the Lightweight Swish-Gated Linear Unit-based Transformer (L-SwiGLU Transformer), optimized for improved performance in resource-constrained scenarios. Furthermore, we perform a comprehensive evaluation of various architectural configurations within Transformers across datasets of varying sizes. This analysis underscores the efficacy of the proposed methodology, illuminating potential avenues for further research and development.
@@
  • We propose L-SwiGLU Transformer, a modified Transformer architecture that enhances computational efficiency and positioning accuracy by replacing the MLP-based feed-forward network (FFN) with a Swish gated linear unit (GLU)-based FFN. Additionally, we replace standard components, such as the class token and normalization layers, with more efficient counterparts. Furthermore, by introducing a global average pooling layer, we demonstrate that positional embeddings can be omitted without adversely affecting positioning accuracy.
@@
The structure of the paper is organized as follows. Section II provides a comprehensive review of the existing literature. The system model is described in detail in Section III. Section IV explains the preprocessing steps and the tokenization methods. The design of transformer architectures for processing distributed sensor signals is discussed in Section V. Extensive evaluation studies and analyses are presented in Section VI. Finally, the paper concludes in Section VII.
@@ II Literature Review
In the literature, indoor positioning is classified into two major categories: geometric-based methods and fingerprinting-based methods [3]. Geometric approaches, such as trilateration and triangulation, use parameters such as AoA, TDoA, ToF, and ToA for positioning. Although these algorithms are well-established and extensively studied, they perform poorly in indoor scenarios due to outlier distortion caused by NLOS and multipath challenges [3].
Conversely, fingerprinting methods involve creating a database by collecting signals from various locations and extracting features from them. This constructed database is then used to predict the location of new signals. Compared to geometric approaches, fingerprinting methods are relatively simple, easily integrated into smart devices, and capable of achieving acceptable accuracy with support from existing wireless infrastructure, which explains their widespread exploration in the literature [3, 8].
Most of the work involving ML-based fingerprinting has utilized shallow ML algorithms like K-nearest neighbors (KNN), Decision Trees (DTs), Support Vector Machines (SVM), and Random Forests. These algorithms are computationally expensive to train on large datasets and their performance tends to degrade. Furthermore, the performance of traditional ML methods will not scale with larger, more complex datasets [9]. With the advent of DL models, researchers have begun exploring algorithms like $\mathsf{MLP}$s, RNNs, LSTMs, and CNNs for fingerprinting. Comprehensive details about deep learning-based fingerprinting methods, including the pros and cons of different methods and various types of inputs used for positioning, are provided in [3, 8].
@@
Numerous DL techniques, including CNNs and LSTMs, have achieved remarkable success in indoor localization [10, 11]. In [12], a fully complex-valued neural network (CVNN) was proposed for positioning in environments with moderate Line-of-Sight (LoS) conditions, demonstrating high accuracy in 2D localization tasks. Building on this work, [13] extended the approach by employing a complex-valued ResNet model tailored for highly NLoS scenarios.
@@
With the recent success of Transformer-based models in various applications, researchers have become increasingly interested in exploring their potential for indoor localization. In [14], the Vision Transformer (ViT) [7] which in encoder only Transformer model was employed for fingerprinting-based indoor localization. The approach involved using a CNN to create patches from the CSI matrix, which were then fed into the Transformer encoder blocks for localization. Similarly, the authors in [15] used ViT for WiFi-assisted indoor localization. They applied Principal Component Analysis (PCA) for input normalization and created RSSI gray images from the normalized RSSI data, achieving a $50^{th}$ percentile error of $1.788$ m. However, the use of PCA makes this approach impractical for real-world applications, as the same transformation cannot be consistently applied during testing, leading to performance degradation. Additionally, the dataset used in this study is not publicly available, making it difficult to assess the proportion of LOS.
@@
In [9], a Transformer encoder-only network is utilized for Received Signal Strength (RSS) based WiFi fingerprinting, where the values from wireless access points (WAPs) serve as inputs. The study focuses on building floor and room prediction tasks using RSS data. However, since only a single RSS value per WAP is available, the model’s ability to extract meaningful spatial or temporal features is inherently limited. Furthermore, the use of PCA as a pre-processing step may not be optimal.
@@
The authors in [16] proposed a hybrid model incorporating CNNs and ViT for localization in Long Range Wide Area Network (LoRaWAN). In this model, CNN is used to learn local features while ViT captures global features. Similarly, in [17], the authors presented a CNN-aided ViT-based indoor localization method. In this approach, the CSI matrix is treated as an image and fed into a CNN, with the resulting features used as patches for the ViT. In the context of indoor localization for MIMO systems, the study in [18] adopts a similar hybrid approach, utilizing a variant of the ViT alongside a CNN for feature extraction.
@@
An approach described in [19] employs a Transformer for signal source localization using the AoA, creating images based on the reference point’s location and arrival angle. However, the process for generating these images is not explained. Additionally, this method utilizes both the encoder and decoder parts of the Transformer, which is unusual since typically only the encoder is used for classification and regression tasks. The work and rationale for this architectural choice are also not provided. ViT have also been investigated for human activity recognition (HAR) using WiFi CSI [20]. In this approach, the raw CSI data is transformed into a CSI spectrogram, which is then used as input to the ViT model for effective HAR.
@@ III System Model
We consider an indoor factory scenario with dense clutter and High sensor height (InF-DH) as in Fig. 1, which illustrates the spatial layout of the 3GPP InF-DH environment, including the arrangement of the sensor nodes connected to a 5G base station, operating at $3.5$ GHz with a signal bandwidth of $100$ MHz. A total of $N_{\rm S}=18$ sensor nodes are uniformly distributed, spaced $20$ m apart, with the perimeter sensor nodes located $10$ m from the walls. The heights of the sensor nodes and devices are fixed at $8$ m and $1.5$ m, respectively. The spatial coordinates of the $r$-th sensor are represented as a $\bm{\psi}_{r}=[x_{\rm s}^{r},y_{\rm s}^{r},z_{\rm s}^{r}]^{\top}\in\mathbb{R}^{3}$. The positions of all $N_{\rm S}$ sensors are then collectively represented by $\bm{P}=[\bm{\psi}_{1}\,\bm{\psi}_{2}\,\cdots\,\bm{\psi}_{N_{\rm S}}]\in\mathbb{R}^{3\times N_{\rm S}}$, where each column corresponds to the coordinates of a single sensor. A single transmit antenna port is assumed for the device. Each sensor node is equipped with a dual-polarized receive antenna, enabling the reception of multi-antenna signals (MAS) through A=2 antenna ports. The RF channels between a device and the sensor nodes are generated using the 3GPP channel model documented in 3GPP Technical Report 38.901 [21].
@@
The 3GPP InF-DH environment features a variety of obstructions, such as small to medium-sized metallic equipment and irregularly shaped items like assembly and assorted machinery. In this paper, we focus on the InF-DH $\{60\%,~{}6~{}\text{m},~{}2~{}\text{m}\}$ configuration where $60$% of the environment is covered by clutters of 6m height and 2m width/length. The LoS probability between a device and a sensor node is 0.008, based on empirical observations from the dataset for the InF-DH scenario [21]. That is, a LoS link between a device and the distributed sensor nodes is very unlikely. To locate a device, the device is configured to transmit the 5G sounding reference symbols (SRS) $\{s_{\rm k}\}_{k=-N/2}^{N/2-1}$, where $N=3264$ is the number of SRS subcarriers, in the uplink to be received by the $18$ sensor nodes. The device transmit power is assumed $23$ dBm. The SRS signals have a comb spacing of $K_{\text{TC}}=2$ subcarriers with $\Delta_{f}=30$ kHz subcarrier spacing and offset between $N_{\text{symb}}=2$ consecutive OFDM symbols. The signal being sampled by each sensor node at a rate of $f_{\text{s}}=4096\cdot\Delta_{\text{f}}=122.88$ MHz is converted into the frequency domain (FD) using an FFT of size $N_{\text{FFT}}=4096$. Assuming $L$ channel taps with complex gains $\{c_{a,l}\}_{l}$ and delays $\{\tau_{l}\}_{l}$ , the received SRS at the $(k+\frac{N}{2})$-th sub-carrier at receive antenna port $a$ is given by
@@
The PDP can be truncated to the first $N_{\text{ts}}=128$ time samples without losing much information about the link between a device and a sensor node. Note that we intentionally keep the processing of the received RF signals to the minimum and leave any signal processing (e.g., filtering, smoothing, or channel estimation) to the DL model.
With $N_{\rm S}=18$ sensor nodes, each device position can be considered to map uniquely to a PDP matrix $\mathbf{\Psi}$ of dimensions $N_{\rm S}\times N_{\textrm{ts}}$, with each row capturing the PDP at a different sensor. We considered three datasets, namely small, medium, and large, sampling 10,$00,0$ 000, $20,000$, and $40,000$ uniformly randomly placed device locations, respectively, and recorded the corresponding PDPs. For the test datasets, we sample another $4,000$ uniformly randomly placed device locations and the corresponding TD PDPs. The PDPs collected from the $N_{\rm S}=18$ sensor nodes are used as inputs to the DL model (to be defined in Section V) for joint processing, which will then output the 2D coordinates of the device.

### v2

@@ Abstract
Indoor localization in challenging non-line-of-sight (NLOS) environments often leads to poor accuracy with traditional approaches. Deep learning (DL) has been applied to tackle these challenges; however, many DL approaches overlook computational complexity, especially for floating-point operations (FLOPs), making them unsuitable for resource-limited devices. Transformer-based models have achieved remarkable success in natural language processing (NLP) and computer vision (CV) tasks, motivating their use in wireless applications. However, their use in indoor localization remains nascent, and directly applying Transformers for indoor localization can be both computationally intensive and exhibit limitations in accuracy. To address these challenges, in this work, we introduce a novel tokenization approach, referred to as Sensor Snapshot Tokenization (SST), which preserves variable-specific representations of power delay profile (PDP) and enhances attention mechanisms by effectively capturing multi-variate correlation. Complementing this, we propose a lightweight Swish-Gated Linear Unit-based Transformer (L-SwiGLU-T) model, designed to reduce computational complexity without compromising localization accuracy. Together, these contributions mitigate the computational burden and dependency on large datasets, making Transformer models more efficient and suitable for resource-constrained scenarios. Experimental results on simulated and real-world datasets demonstrate that SST and L-SwiGLU-T achieve substantial accuracy and efficiency gains, outperforming larger Transformer and CNN baselines by over 40% while using significantly fewer FLOPs and training samples.
@@ I Introduction
The rapid advancement of wireless communication technologies, fueled by the deployment of 5G and the emerging vision of 6G, is driving transformative innovations in location-based services. Applications such as vehicle-to-everything (V2X) communication, urban air mobility (UAM), extended reality (XR), near-real-time robotic operations, and Smart X systems [lee2023towards, nguyen2019localization] are at the forefront of this technological evolution. These advancements require the large-scale integration of Internet of Things (IoT) infrastructures, thereby intensifying the demand for accurate and efficient localization methods with reduced computational complexity to ensure the stringent quality of service (QoS) necessary for these cutting-edge applications. Indoor localization is particularly critical for applications such as navigation, safety and rescue operations, and resource optimization for customized user services. However, most indoor positioning devices remain constrained by limited computing power, highlighting the need for localization methods that are both efficient and accurate while supporting the demands of next-generation wireless systems.
@@
Global Navigation Satellite Systems (GNSS), commonly used for outdoor localization, face significant challenges in indoor environments due to severe channel conditions such as shadowing, fading, and noise, as well as the high probability of non-line-of-sight (NLOS) situations. Conventional wireless localization approaches often rely on signal-level metrics such as Time of Arrival (ToA), Time Difference of Arrival (TDoA), Time of Flight (ToF), Angle of Arrival (AoA), and Received Signal Strength Indicator (RSSI) [singh2021machine]. These metrics are commonly used in geometric or statistical localization methods, which perform poorly in NLOS-rich indoor environments such as factories, hospitals, and shopping malls due to their reliance on complex, environment-dependent empirical models.
@@
Fingerprinting [shang2022overview] is a widely used indoor localization approach that leverages signal measurements such as Received Signal Strength Indicator (RSSI) or Channel State Information (CSI). However, traditional machine-learning methods struggle with environmental variability and fail to capture complex spatial–temporal dependencies. Deep learning (DL) models—such as Convolutional Neural Networks (CNNs) and Multi-Layer Perceptrons (MLPs) [singh2021machine] learn more discriminative features and offer improved robustness, but they also increase system complexity and computational cost, hindering real-time implementation and scalability. Therefore, this work aims to achieve a balance between accuracy and efficiency, enabling adaptable and lightweight fingerprinting-based localization.
@@
The Transformer model [vaswani2017attention] offers several advantages over traditional sequence processing techniques such as Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks, including parallelization capability, reduced training time, and improved handling of long-range dependencies [khan2022transformers]. Transformers, despite their proven success in fields such as natural language processing (NLP), computer vision (CV), and machine translation [vaswani2017attention, dosovitskiy2020image], have not been extensively used for indoor localization. This underutilization is largely due to the lack of inductive biases such as locality and translation invariance, which are naturally present in CNNs but absent in Transformers, and due to the Transformer’s dependence on large datasets to capture both local and global features [dosovitskiy2020image]. These challenges are further amplified by the way inputs are tokenized, which can impact the model’s training effectiveness.
@@
In highly NLOS indoor wireless environments, multi-antenna systems face significant challenges due to multipath propagation, where signals undergo reflection, diffraction, and scattering from walls, furniture, and other obstacles. This complex propagation leads to signal superposition, causing phase shifts, fading, and interference, which further complicate the resolution of spatial and temporal variations in such environments. Moreover, in the context of indoor localization, collecting large-scale localization datasets is costly and time-consuming, while existing infrastructures lack the computational capacity to test models requiring billions of FLOPs. These challenges underscore the need for more efficient approaches that reduce the dependence of Transformer-based models on large datasets while simultaneously lowering their computational complexity, enabling practical deployment in indoor wireless environments.
@@
One critical yet often overlooked aspect of Transformers, particularly within the wireless communication domain, is tokenization. Conventional tokenization methods typically aggregate multiple variables, such as delayed events and physical measurements, into a single representation. This aggregation can obscure variable-specific features, leading to less informative attention maps and degrading model performance. Consequently, the model’s dependency on extensive datasets increases. Therefore, in this paper, we investigate the Transformer architecture and enhance its learning capabilities by proposing a sophisticated and physically interpretable tokenization technique referred to as Sensor Snapshot Tokenization (SST) that leverages the characteristics of wireless communication systems, specifically addressing inherent channel independence and enabling the Transformer to effectively learn the multivariate correlations using the tokens form multi-antenna radio system. SST produces variate-centric representations with more meaningful attention maps, reducing the need for large models and extensive training data. To further enhance efficiency and reduce computational complexity, we introduce a lightweight Transformer variant, termed the Lightweight Swish-Gated Linear Unit Transformer (L-SwiGLU-T). Building upon the Vanilla Transformer, L-SwiGLU-T incorporates selective normalization, compact feed-forward design, and an efficient prediction mechanism, collectively improving stability and reducing FLOPs.
@@
  • We propose L-SwiGLU-T, a modified Transformer architecture that enhances computational efficiency and positioning accuracy by replacing the MLP-based feed-forward network (FFN) with a Swish gated linear unit (GLU)-based FFN. Additionally, we replace standard components, such as the class token and normalization layers, with more efficient counterparts. Furthermore, by introducing a global average pooling (GAP) layer, we demonstrate that positional embeddings can be omitted without adversely affecting positioning accuracy.
  • We added a comparison of the proposed model against two recent lightweight Transformer architectures, and validated its performance on a real-world dataset.
@@
Our preliminary study in [11162366] introduced the SST method and demonstrated its initial performance improvements using vanilla Transformer. In contrast, this paper extends that work by introducing a lightweight Transformer architecture (L-SwiGLU-T), evaluating across multiple datasets and model sizes, and validating performance on real-world measurements.
The paper is organized as follows. Section II provides a comprehensive review of the existing literature. The system model is described in detail in Section III. Section IV explains the preprocessing steps and the tokenization methods. The design of transformer architectures for processing distributed sensor signals is discussed in Section V. Extensive evaluation studies and analyses are presented in Section VI. Finally, the paper concludes in Section VII.
@@ II Literature Review
In the literature, indoor positioning is classified into two major categories: geometric-based methods and fingerprinting-based methods [singh2021machine]. Geometric approaches, such as trilateration and triangulation, use parameters such as AoA, TDoA, ToF, and ToA for positioning. Although these algorithms are well-established and extensively studied, they perform poorly in indoor scenarios due to outlier distortion caused by NLOS and multipath challenges [singh2021machine].
Conversely, fingerprinting methods construct a database of signal features collected from known locations to predict the position of new measurements. Compared with geometric approaches, they are simpler, easily integrated into smart devices, and can achieve good accuracy using existing wireless infrastructure [singh2021machine, alhomayani2020deep]. Early ML-based fingerprinting primarily relied on shallow algorithms such as K-nearest neighbors (KNN), Decision Trees (DTs), Support Vector Machines (SVMs), and Random Forests, which are costly to train and perform poorly on large or complex datasets [nasir2024hytra]. With the advent of deep learning, models such as Multilayer Perceptrons (MLPs), RNNs, LSTMs, and CNNs have been explored, offering improved feature learning and generalization. A comprehensive overview of DL-based fingerprinting, including model comparisons and input modalities, is provided in [singh2021machine, alhomayani2020deep].
@@
Numerous DL techniques, including CNNs and LSTMs, have achieved remarkable success in indoor localization [hoang2020cnn, gao2022toward]. In [cheng2024], a fully complex-valued neural network (CVNN) was proposed for positioning in environments with moderate Line-of-Sight (LoS) conditions, demonstrating high accuracy in 2D localization tasks. Building on this work, [khamesi2024] extended the approach by employing a complex-valued ResNet model tailored for highly NLoS scenarios.
@@
With the growing success of Transformer architectures, several studies have explored their use in indoor localization. In [li2023lot], a Vision Transformer (ViT) [dosovitskiy2020image], an encoder-only Transformer, was applied to fingerprinting-based localization, where a CNN generated CSI patches that were processed by Transformer encoders. Similarly, [zhou2024vtil] employed ViT for WiFi-assisted localization, using Principal Component Analysis (PCA) for input normalization and converting RSSI data into grayscale images, achieving a 50th-percentile error of $1.788$ m. However, PCA-based preprocessing is impractical for deployment since the transformation cannot be consistently applied during inference, leading to degraded performance. Moreover, the dataset used in [zhou2024vtil] is not publicly available, limiting reproducibility and evaluation under different LoS/NLoS conditions.
@@
In [nasir2024hytra], a Transformer encoder-only network is utilized for Received Signal Strength (RSS) based WiFi fingerprinting, where the values from wireless access points (WAPs) serve as inputs. The study focuses on building floor and room prediction tasks using RSS data. However, since only a single RSS value per WAP is available, the model’s ability to extract meaningful spatial or temporal features is inherently limited. Furthermore, the use of PCA as a pre-processing step may not be optimal.
@@
The authors in [lutakamale2024hybrid] proposed a hybrid model incorporating CNNs and ViT for localization in Long Range Wide Area Network (LoRaWAN). In this model, CNN is used to learn local features while ViT captures global features. Similarly, in [prasad2024vision], the authors presented a CNN-aided ViT-based indoor localization method. In this approach, the CSI matrix is treated as an image and fed into a CNN, with the resulting features used as patches for the ViT. In the context of indoor localization for MIMO systems, the study in [xu2024swin] adopts a similar hybrid approach, utilizing a variant of the ViT alongside a CNN for feature extraction.
@@
An approach described in [pan2022transformer] employs a Transformer for signal source localization using the AoA, creating images based on the reference point’s location and arrival angle. However, the process for generating these images is not explained. Additionally, this method utilizes both the encoder and decoder parts of the Transformer, which is unusual since typically only the encoder is used for classification and regression tasks. The work and rationale for this architectural choice are also not provided. ViT have also been investigated for human activity recognition (HAR) using WiFi CSI [luo2024vision]. In this approach, the raw CSI data is transformed into a CSI spectrogram, which is then used as input to the ViT model for effective HAR.
@@ III System Model
We consider an indoor factory scenario with dense clutter and High sensor height (InF-DH) as in Fig. 1, which illustrates the spatial layout of the 3GPP InF-DH environment, including the arrangement of the sensor nodes connected to a 5G base station, operating at $3.5$ GHz with a signal bandwidth of $100$ MHz. A total of $N_{\rm S}=18$ sensor nodes are uniformly distributed, spaced $20$ m apart, with the perimeter sensor nodes located $10$ m from the walls. The heights of the sensor nodes and devices are fixed at $8$ m and $1.5$ m, respectively. The spatial coordinates of the $r$-th sensor are represented as a $\bm{\psi}_{r}=[x_{\rm s}^{r},y_{\rm s}^{r},z_{\rm s}^{r}]^{\top}\in\mathbb{R}^{3}$. The positions of all $N_{\rm S}$ sensors are then collectively represented by $\bm{P}=[\bm{\psi}_{1}\,\bm{\psi}_{2}\,\cdots\,\bm{\psi}_{N_{\rm S}}]\in\mathbb{R}^{3\times N_{\rm S}}$, where each column corresponds to the coordinates of a single sensor. A single transmit antenna port is assumed for the device. Each sensor node is equipped with a dual-polarized receive antenna, enabling the reception of multi-antenna signals (MAS) through A=2 antenna ports. The RF channels between a device and the sensor nodes are generated using the 3GPP channel model documented in 3GPP Technical Report 38.901 [3GPP_TR_38_901].
@@
The 3GPP InF-DH environment features a variety of obstructions, such as small to medium-sized metallic equipment and irregularly shaped items like assembly and assorted machinery. In this paper, we focus on the InF-DH $\{60\%,~6~\text{m},~2~\text{m}\}$ configuration where 6060% of the environment is covered by clutters of 6m height and 2m width/length. The LoS probability between a device and a sensor node is 0.008, based on empirical observations from the dataset for the InF-DH scenario [3GPP_TR_38_901]. That is, a LoS link between a device and the distributed sensor nodes is very unlikely. To locate a device, the device is configured to transmit the 5G sounding reference symbols (SRS) $\{s_{\rm k}\}_{k=-N/2}^{N/2-1}$, where $N=3264$ is the number of SRS subcarriers, in the uplink to be received by the $18$ sensor nodes. The device transmit power is assumed $23$ dBm. The SRS signals have a comb spacing of $K_{\text{TC}}=2$ subcarriers with $\Delta_{f}=30$ kHz subcarrier spacing and offset between $N_{\text{symb}}=2$ consecutive OFDM symbols. The signal being sampled by each sensor node at a rate of $f_{\text{s}}=4096\cdot\Delta_{\text{f}}=122.88$ MHz is converted into the frequency domain (FD) using an FFT of size $N_{\text{FFT}}=4096$. Assuming $L$ channel taps with complex gains $\{c_{a,l}\}_{l}$ and delays $\{\tau_{l}\}_{l}$ , the received SRS at the $(k+\frac{N}{2})$-th sub-carrier at receive antenna port $a$ is given by
@@
The PDP can be truncated to the first $N_{\text{ts}}=128$ time samples without losing much information about the link between a device and a sensor node. Note that we intentionally keep the processing of the received RF signals to the minimum and leave any signal processing (e.g., filtering, smoothing, or channel estimation) to the DL model. Unlike RSS or AoA-based features, which suffer from instability and degraded accuracy in NLOS conditions, the proposed approach utilizes the PDP as the input representation. While CSI provides richer frequency-domain information, it is complex-valued and high-dimensional, leading to substantially higher computational overhead. In contrast, the PDP offers a real-valued, compact, and stable time-domain representation of multipath propagation, making it well-suited for the lightweight and resource-efficient design.
With $N_{\rm S}=18$ sensor nodes, each device position can be considered to map uniquely to a PDP matrix $\mathbf{\Psi}$ of dimensions $N_{\rm S}\times N_{\textrm{ts}}$, with each row capturing the PDP at a different sensor. We considered three datasets ($\mathcal{D}$), namely small, medium, and large, sampling 10,0000,000, $20,000$, and $40,000$ uniformly randomly placed device locations, respectively, and recorded the corresponding PDPs. For the test datasets, we sample another $4,000$ uniformly randomly placed device locations and the corresponding TD PDPs. The PDPs collected from the $N_{\rm S}=18$ sensor nodes are used as inputs to the DL model (to be defined in Section V) for joint processing, which will then output the 2D coordinates of the device.
III-A Problem Formulation
We aim to learn a parametric model $f_{\bm{\Theta}}$ that estimates the 2D device location $\hat{\mathbf{y}}\in\mathbb{R}^{2}$ from the input PDPs $\bm{\Psi}\in\mathbb{R}^{N_{\mathrm{S}}\times N_{\mathrm{ts}}}$. The Transformer model parameters are denoted by $\bm{\Theta}$, and the predicted location is given by $\hat{\mathbf{y}}=f_{\bm{\Theta}}(\bm{\Psi})$, with the ground-truth location $\mathbf{y}\in\mathbb{R}^{2}$.
The learning objective is to minimize the expected 2D localization error, measured by the $L_{2}$ norm, while enforcing a computational constraint and ensuring robustness under data-limited conditions. This can be expressed as the following regularized optimization problem:
$\min_{\bm{\Theta}}$ $\mathbb{E}_{(\bm{\Psi},\mathbf{y})\sim\mathcal{D}}\!\left[\left\|f_{\bm{\Theta}}(\bm{\Psi})-\mathbf{y}\right\|_{2}^{2}\right]+\mathcal{R}_{\mathcal{D}}\!\left(f_{\bm{\Theta}}\right)$ (5)
s.t. $\mathrm{FLOPs}\!\left(f_{\bm{\Theta}}\right)\leq\mathcal{B}$
The first term in (5) represents the mean-squared localization loss, which directly drives the optimization of the model parameters $\bm{\Theta}$. The second term, $\mathcal{R}_{\mathcal{D}}\!\left(f_{\bm{\Theta}}\right)$, acts as a data-efficiency regularizer that enhances model robustness under limited training data. This robustness is further promoted through the proposed tokenization strategy, proposed Transformer-based model, and data augmentation, which together reduce dependence on large datasets and enable effective learning from diverse PDP realizations. The constraint on $\mathrm{FLOPs}\left(f_{\bm{\Theta}}\right)$ enforces a computational budget $\mathcal{B}$, ensuring that the total computational cost of the model does not exceed this limit. In this work, three model configurations are defined: small, medium, and large, corresponding to computational budgets of approximately $\mathcal{B}=4.5$M, 16.516.5M, and 63.563.5M FLOPs, respectively. The proposed architecture adheres to this constraint through a series of design optimizations, such as removing positional embeddings and the class token, adopting GAP, and carefully selecting the number of Transformer layers as well as the embedding and hidden dimensions.